OpenStudy (anonymous):

How do you solve this (cos(θ)-1)/(1-sec(θ))=(cos(θ)+1)/(1+sec(θ))

OpenStudy (anonymous):

True?

OpenStudy (anonymous):

is this trigonometry?

OpenStudy (anonymous):

yes, it's trig identities but I don't know how to prove it's true.

OpenStudy (anonymous):

It is true trust me

OpenStudy (anonymous):

okay, thank you. can you show me how to solve it?

OpenStudy (tkhunny):

Cosines and secants? Isn't it just dying for you to change the secants to cosines and then just see what happens?

OpenStudy (anonymous):

but when you change them to cosines won't it be (cos(θ)-1)/(1-(1)/(cos(θ))=(cos(θ)+1)/(1+(1)/(cos(θ)))

OpenStudy (tkhunny):

Sure. Then a little algebra...

OpenStudy (anonymous):

okay thanks :) I'll try that

OpenStudy (anonymous):

LHS

OpenStudy (anonymous):

\[\frac{ \cos \theta -1 }{ 1- \sec \theta }= \frac{ \cos \theta -1 }{ 1- \frac{ 1 }{ \cos \theta } }\] \[=\frac{ \cos \theta -1 }{ \frac{ \cos \theta -1 }{ \cos \theta } }= \cos \theta\]

OpenStudy (anonymous):

RHS do the same, and got the same answer = cos theta ---> LHS = RHS

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Sorry, My15girl, I wasn't ignoring you. I was trying to figure out your problem when the system threw me off-line. It's just as well, though. It appears that Hoa has a much better grasp of the relationship between secants and cosines than I currently do, and aided you much better than I could have. :-)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Sorry I didn't answer you earlier, qweqwe123123123123111 after Hoa helped me I had to to go and logged off, :( I said I'd come back and look at it later but, I was having trouble logging back in, because I couldn't verify my email. But I did in the end. So glad too. Anyway I really appreciate you trying to help me. It was very kind. Thank you so much. The effort means a lot :)